# Predicting the single-site and multi-site event discrimination power of dual-phase time projection chambers

### v2

@@ 3.2 Maximum likelihood reconstruction
where the “hat” symbol indicates the best-fit parameter values. $q$ measures the improvement of the fit quality with a MS2 hypothesis over a SS hypothesis. It is also a statistic that is invariant under change of variables, and can indicate MLE’s capability in resolving MS2 events.
@@
Because the MS2 fit includes additional parameters and should, on average, obtain a larger likelihood value, the likelihood ratio should be bounded between 0 and 1. Therefore, $q$ should be a positive number and the factor of 2 is added so the distribution of $q$ for SS events follow a $\chi^{2}$ distribution (Wilk’s theorem). The $q$ distribution for SS events, and MS2 events with inter-vertex distances of 3 and 6 cm are shown in Fig. 11 (left). For SS events, the two likelihood values should be close to each other, yielding $q\sim$0, while for resolvable MS2 events the MS2 fit should return a much larger likelihood, causing $q>0$. The MLE’s ability to distinguish MS and SS events resides in the deviation of the $q$ distribution from 0 for MS2 events.
@@
To obtain a characteristic MS2 resolution that can be compared to the benchmark theoretical prediction, we choose the inter-vertex separation value $d$ that leads to 16% MS2 contamination when a cut on $q$ value accepts 50% of SS events. For 500 detected photons the characteristic MS2 resolution is estimated to be 1.7cm, which is slightly larger than the theoretical limit of 1.2 cm. The characteristic resolution dependence on the number of detected photons (assuming 1-1 splitting between two vertices) is shown in Fig. 11 (right), in comparison with that predicted by FI and CRLB (Fig. 6). The MLE is observed to outperform the PCA for most of the scenarios tested.
3.3 Classification using Convolutional Neural Network
@@
We simulate SS and MS2 events in the central region of a large TPC instrumented with 2" PMTs. SS event positions are drawn uniformly from within a central disk in the TPC. For MS2 events, the distance from the first vertex to the second, and the angle defining the direction of the second vertex are both sampled from uniform distributions. While separation between vertices in physical events, such as those produced by Compton scatters of gamma rays, are often exponentially distributed, we use the uniform distribution to ensure training data span a sufficient range of separations. As done in studies presented in early sections, the mean number of detected photons at each sensor is sampled from a Poisson distribution with a mean determined from an analytical LRF and the radial distance to the event vertex.
@@
The sensor hit pattern in each simulated event produces an image with which we can train or test the CNN. First, the sensor nearest the photon centroid is chosen as the center of the image. Then, a rectangular image is constructed by taking a fixed number of rows and columns along a selected direction. Due to the six-fold rotational symmetry of the sensor array arising from the hexagonal packing of the circular sensors, we choose the direction in which the second moment of the light detected is maximal. This ensures that the variance in light collection is predominantly along the horizontal axis of each image, reducing the number of patterns the CNN must learn. The resulting rows of sensors are then mapped to a rectangular grid of pixels by interpolating the pixel value in the alternate columns. To ensure that the CNN does not use the total energy of an event, and leverages only the spatial information, the number of detected counts at each pixel is then normalized by the sum over the full image. This process is illustrated in Fig. 12, which shows a sample sensor hit pattern in the left and the resulting image used to train the CNN in the right.
Architecture and hyperparameter selections are key steps in the design of a ML-based classifier. We use a ResNet-14 [24] architecture with 8, 16, and 32 filters at each convolutional layer in the ResNet blocks, based on the size of our training datasets and classification task. While a simpler CNN architecture could achieve similar performance, the ResNet architecture is easily scalable to larger training datasets for more complex tasks such as classification of events with any number of vertices and energy partitioning between sites. Similar results were obtained with a small-scale VGG-style architecture [25]. The ResNet-14 architecture has $\sim\!4.5\times 10^{4}$ parameters, and the training data consists of $10^{5}$ images each for the SS and MS categories. The choices ensure at least as many training samples as trainable parameters. The CNN is trained using a binary cross-entropy loss function to categorize events as SS or MS. Learning rate scheduling and early stopping were implemented to maximize performance of the classifier on the validation data.
Figure 13: Left: MS rejection efficiency as a function of the vertex separation and photon budget for double vertex events with equal photon splitting. These curves assume 50% signal acceptance. Right: Benchmark MS2 resolution as a function of the number of photons detected using the CNN (solid line) compared to the CRLB (dotted line). We assume the use of 2" PMTs in this study.
@@ 4.1 Background rejection in neutrinoless double beta decay experiments
A key scientific application of SS/MS discrimination in liquid xenon detectors is the search for $0\nu\beta\beta$ decays of 136Xe, for which the signal is the simultaneous emission of two $\beta$ electrons from the same nucleus with total energy summing to $Q=2.457$ MeV. At these energies, electrons have a typical track length of $\mathcal{O}$(1 mm) in liquid xenon, and therefore the $0\nu\beta\beta$ signals will primarily appear as SS events, with $\sim$10% that appear as MS events due to the emission of Bremsstrahlung photons from the electron tracks. The dominant background near the $0\nu\beta\beta$ $Q$-value in TPC-based experiments is the 2.447 MeV gamma ray emitted by the $\beta$ decay of 214Bi, which comes from the primordial 238U decay chain and is present at trace levels in detector materials [16, 15]. These gammas have a percent-level probability of direct photoabsorption in xenon, which would deposit all the energy at a single site and would not be easily differentiated from $0\nu\beta\beta$ decays. However, most of the time the gamma ray will lose a fraction of its energy by Compton scattering at least once before being fully absorbed, and therefore could be identified as MS events. The most important subset of these events are two-scatter events, in which the first scatter deposits the maximum allowed Compton energy, leaving only $\sim$250 keV in the scattered photon. Since the mean free path of photons decreases with energy, these events will have a smaller average separation between vertices than the typical gamma event, and thus are the primary background pathology.
@@
Figure 14: The distributions of fractional energy deposits by 214Bi gamma-ray interactions in liquid xenon with different vertex multiplicities (left) and the cumulative position distributions between vertices in double-scatter events before and after a $z$ cut (right). The data is obtained from a Monte Carlo simulation with Geant4.
At a nominal electric field of 200V/cm, a 1.23MeV electron can produce $\sim$60,000 ionization electrons in liquid xenon, leading to $\sim$120,000 total electrons in a single 136Xe $0\nu\beta\beta$ event if cross-track electron-ion recombination [27] is not considered. If we further assume an ionization signal gain of 20 photoelectrons/e- with half of them detected by sensors in the top, a $0\nu\beta\beta$ event could contain $\sim 1.2\times 10^{6}$ detectable photoelectrons for xyx-y position reconstruction, of which up to 300,000 could be detected by a single 2" PMT in our default detector configuration. Here we choose a relatively low electron gain to mitigate PMT saturation issues while allowing for single ionization electrons to be identified. However, such high levels of light output in a short time window ($\sim 1\mu s$) can still cause the PMTs to saturate. This issue may be further mitigated by reading out the PMT signals from one of the last dynodes instead from the anode, or by optimizing the tapering ratio of the PMT voltage divider. It has been demonstrated that such saturation effects can also be partially corrected in analysis by focusing on the rising edges of the pulses [28]. Saturation leads to an additional uncertainty in the number of photons detected in the affected channel, which could degrade the performance of the SS/MS reconstruction. The following study ignores such photosensor saturation effects, assuming that either software or hardware solutions can be found to mitigate their impact in the experiment.
Sec. 2.3 estimates a theoretical limit of 7.3 mm for the benchmark MS2 rejection resolution with $1.2\times 10^{6}$ detected photons, assuming the use of 2" PMTs and a 1-9 photon splitting between two vertices. This benchmark resolution corresponds to a rejection of 84% MS2 backgrounds separated by this distance while accepting 50% SS signals. The background rejection power deteriorates for events with smaller separations between vertices or with more imbalanced energy distribution. Full calculations of the $0\nu\beta\beta$ signal acceptance and 214Bi background rejection power in our theoretical framework are not straightforward, so we use the MLE and CNN classifier to approximate the TPC’s capability in this application. As demonstrated in Sec. 3, these methods can approach the theoretical CRLB limits in idealized scenarios. Similar to earlier studies, we use an analytical light response function for both sensor hit pattern simulation and reconstruction, and thus ignore uncertainties beyond statistical fluctuations.
@@ 4.2 Experiment optimization
The CRLB derived from the FI matrix is an estimate of the best reconstruction accuracy for parameters of interest with an unbiased estimator. On the one hand, this limit may not be achieved in experiments, especially when simplifications are made in modeling the system behavior (forward model) during sensitivity predictions. Under favorable conditions, however, the limits could be reasonably approached; for example, the SS position resolution achieved by the LUX experiment is within 20% from the predictions made in Appendix. B. Further, the comparison of performance for different design options will remain valid if they share similar uncertainty sources. As a result, even when the predicted performance limit could not be reached, FI remains a useful optimization tool.
@@
Figure 16: Predicted SS (left) and MS2 (right) reconstruction resolutions for 2" photosensors (red lines) and 6 mm photosensors (blue lines) with different distances between the electroluminescence region and the photosensor array (denoted as $h$): 6 cm (solid lines), 1 cm (dashed lines), 0.5 cm (dotted lines). For MS2 events a 1-1 energy splitting between the two sites is assumed.
In our TPC model, the separation between MS event vertices is governed by physical laws and cannot be increased, while the observed spread of the light response function in a TPC can be strongly shaped by choices of detector configuration. Smaller sensor pixels can lead to sharper light response functions and better position resolution, as shown in Figs. 2 and 6. For sensor sizes comparable to, or smaller than the distance between the photon source and the sensor plane, the gain from reducing sensor size diminishes. Therefore, the benefits of small sensors can only be realized by choosing a small physical separation between the light source and the sensor plane. This is demonstrated in Fig. 16, where we compare the predicted SS (left) and MS2 (right, assuming 1-1 photon splitting between two vertices) resolutions for varying distances between the electroluminescence region and the photosensor plane, assuming the use of 2" sensors (red lines) and 6 mm sensors (blue lines). Compared to the baseline model of 1 cm sensor separation (dashed lines), a 0.5 cm separation  [^3] For the 0.5 cm sensor separation study, the electroluminescence region thickness is also reduced to 0.5 cm. (dotted lines) improves the resolution for 6 mm sensors by approximately a factor of 2; the results for 2" sensors remain largely unchanged because the photon distribution is dominated by the large size of the sensor in these two cases. When the sensor separation from the electroluminescence region is increased from 1 cm to 6 cm (solid lines), which was the configuration of the LUX detector, both SS and MS resolutions substantially deteriorate. In this configuration, the 2" sensor results and the 6 mm sensor results are nearly identical, because the photon distributions in both scenarios are dominated by the large light source-sensor separation.
@@ Appendix A Fisher Information for independent Poisson distributions
Appendix B Comparison with LUX experimental result
The LUX experiment operated a dual-phase xenon TPC to search for dark matter interactions between 2012 and 2016 [5]. The LUX TPC used 61 2" PMTs in the xenon gas phase to collect xenon electroluminescence light. The TPC was 50 cm in diameter with a 0.5 cm electroluminescence region between the liquid surface and the TPC anode, and the PMT array was placed 6 cm above the anode. The LUX detector LRF is reported in [14] and its radial component is shown in Fig. 17 (left), in comparison with our simulated LRF (right) in a similar configuration (6 cm between the anode and the PMT array, and a 1 cm electroluminescence region). Despite slightly different detector configurations, the two LRF functions are nearly identical.
Figure 17: Left - radial component of the LUX LRF; both data and fit functions are taken from [14]. Right - simulated LRF in our work, assuming close-packed 2" PMTs and a 1 cm electroluminescence region which is 6 cm away from the PMT array; fit function follow the form of Eq. 2.1.
Using the functional form of the LUX LRF given in [14], we carried out a similar study to predict the theoretical SS event reconstruction resolution in the LUX detector using FI. Although LUX used 2" PMTs (nominal diameter), the distance between neighboring PMTs (center to center) was 6.125 cm. We create a mock-up LUX detector model consisting of 61 PMTs on a hexagonal grid of 6.125 cm pitch in a 50 cm array. Then we simulate SS events near the center of the detector so the polar component of the LUX LRF is relatively small. The predicted SS reconstruction resolution curves for events in the center of the TPC (red solid line), between 2 central PMTs (red dashed line) and at the center of 3 central PMTs (red dotted line), which were referred to as positions A, B, and C, respectively in Fig. 3, are plotted on top of experimental result reported by LUX in Fig. 18. Because of the large light source-PMT array distance, the SS reconstruction performances for these 3 different event positions are similar, which are marginally better than those achieved in the LUX experiment.
Figure 18: Predicted SS event reconstruction resolutions as a function of detected photons for 3 event positions (see text), plotted on top of experimental result from the LUX experiment [14].

### v3

@@ 3.2.2 Likelihood ratio analysis
where the “hat” symbol indicates the best-fit parameter values. $q$ measures the improvement of the fit quality with a MS2 hypothesis over a SS hypothesis. It is also a statistic that is invariant under change of variables, and can indicate MLE’s capability in resolving MS2 events. More importantly, despite its construction, this method retains some power for SS/MS event discrimination even when the number of vertices in an event is not exactly 2, as demonstrated in Sec. 4.1.
@@
Because the MS2 fit includes additional parameters and should, on average, obtain a larger likelihood value, the likelihood ratio should be bounded between 0 and 1. Therefore, $q$ should be a positive number and the factor of 2 is added so the distribution of $q$ for SS events follow a $\chi^{2}$ distribution (Wilk’s theorem). The $q$ distribution for SS events, and MS2 events with inter-vertex distances of 3 and 6 cm are shown in Fig. 12 (left). Here we conveniently use the PCA-estimated principal direction and variance to initialize the MLE fit. For SS events, the two likelihood values should be close to each other, yielding $q\sim$0, while for resolvable MS2 events the MS2 fit should return a much larger likelihood, causing $q>0$. The MLE’s ability to distinguish MS and SS events resides in the deviation of the $q$ distribution from 0 for MS2 events.
@@
To obtain a characteristic MS2 resolution that can be compared to the benchmark theoretical prediction, we choose the inter-vertex separation value $d$ that leads to 16% MS2 contamination when a cut on $q$ value accepts 50% of SS events. For 500 detected photons the characteristic MS2 resolution is estimated to be 1.7 cm, which is slightly larger than the theoretical limit of 1.2 cm. The characteristic resolution dependence on the number of detected photons (assuming 1:1 splitting between two vertices) is shown in Fig. 12 (right), in comparison with that predicted by FI and CRLB (Fig. 6). The MLE is observed to outperform the PCA for most of the scenarios tested.
3.2.3 Convolutional neural network
@@
When simulating events in a large TPC instrumented with 2" PMTs, SS event positions are drawn uniformly from within a central disk. For MS2 events, the distance from the first vertex to the second, and the angle defining the direction of the second vertex are both sampled from uniform distributions. While separation between vertices in physical events, such as those produced by Compton scatters of gamma rays, are often exponentially distributed, we use the uniform distribution to ensure training data span a sufficient range of separations. As done in studies presented early in this section, the mean number of detected photons at each sensor is sampled from a Poisson distribution with a mean determined from an analytical LRF and the radial distance to the event vertex.
@@
The sensor hit pattern in each simulated event produces an image with which we can train or test the CNN. First, the sensor nearest the photon centroid is chosen as the center of the image. Then, a rectangular image is constructed by taking a fixed number of rows and columns along a selected direction. Due to the six-fold rotational symmetry of the sensor array arising from the hexagonal packing of the circular sensors, we choose the direction in which the second moment of the light detected is maximal. This ensures that the variance in light collection is predominantly along the horizontal axis of each image, reducing the number of patterns the CNN must learn. The resulting rows of sensors are then mapped to a rectangular grid of pixels by interpolating the pixel value in the alternate columns. To ensure that the CNN does not use the total energy of an event, and leverages only the spatial information, the number of detected counts at each pixel is then normalized by the sum over the full image. This process is illustrated in Fig. 13, which shows a sample sensor hit pattern on the left and the resulting image used to train the CNN on the right.
The obtained 2D images are then processed using a ResNet-14 [25] architecture with 8, 16, and 32 filters at each convolutional layer. While the ResNet architecture can be scaled to larger training datasets for more complex tasks, a simpler CNN architecture such as a small-scale VGG [26] was found to yield a similar performance in our study. The CNN has $\sim\!4.5\times 10^{4}$ parameters, and the training data consists of $10^{5}$ images each for the SS and MS categories. The choices ensure at least as many training samples as trainable parameters. The CNN is trained using a binary cross-entropy loss function to categorize events as SS or MS. Learning rate scheduling and early stopping were implemented to maximize performance of the classifier on the validation data.
Figure 14: Left - MS rejection efficiency as a function of the vertex separation and photon budget for double vertex events with equal photon splitting. These curves assume 50% signal acceptance. Right - benchmark MS2 resolution as a function of the number of photons detected using the CNN (solid line) compared to the CRLB (dotted line). We assume the use of 2" PMTs in this study.
@@ 4.1 Background rejection in neutrinoless double beta decay experiments
A key scientific application of SS/MS discrimination in liquid xenon detectors is the search for $0\nu\beta\beta$ decays of 136Xe, for which the signal is the simultaneous emission of two $\beta$ electrons from the same nucleus with total energy summing to $Q=2.457$ MeV. At these energies, electrons have a typical track length of $\mathcal{O}$(1 mm) in liquid xenon, and therefore the $0\nu\beta\beta$ signals will primarily appear as SS events, with $\sim$10% that appear as MS events due to the emission of Bremsstrahlung photons from the electron tracks. The dominant background near the $0\nu\beta\beta$ $Q$-value in TPC-based experiments is the 2.447 MeV gamma ray emitted in the $\beta$ decay of 214Bi, which comes from the primordial 238U decay chain and is present at trace levels in detector materials [15, 14]. These gammas have a percent-level probability of direct photoabsorption in xenon, which would deposit all the energy at a single site and would not be easily differentiated from $0\nu\beta\beta$ decays. However, most of the time the gamma ray will lose a fraction of its energy by Compton scattering at least once before being fully absorbed, and therefore could be identified as MS events. The most important subset of these events are two-scatter events, in which the first scatter deposits the maximum allowed Compton energy, leaving only $\sim$250 keV in the scattered gamma to be photoabsorbed. Since the mean free path of gamma rays decreases with energy, these MS2 events will have a small average separation between vertices, and thus are the primary background pathology.
@@
Figure 15: left: the energy fraction contained in the largest cluster from 214Bi gamma-ray interactions in liquid xenon with one, two, and more vertices; the green peak around 0.6 corresponds to pair production by gamma rays. right: the cumulative position distributions between vertices in double-scatter events before and after a 3 mm $z$ cut. The data is obtained from a Monte Carlo simulation with Geant4.
At a nominal electric field of 200 V/cm, a 1.23 MeV electron can produce $\sim$60,000 ionization electrons in liquid xenon, leading to $\sim$120,000 total electrons in a single 136Xe $0\nu\beta\beta$ event if cross-track electron-ion recombination [28] is not considered. If we further assume an ionization signal gain of 20 photoelectrons/e- with half of them detected by sensors in the top, a $0\nu\beta\beta$ event could contain $\sim 1.2\times 10^{6}$ detectable photoelectrons for xyx-y position reconstruction, of which up to 300,000 could be detected by a single 2" PMT in our default detector configuration. Here we choose a relatively low electron gain to mitigate PMT saturation issues while allowing for single ionization electrons to be identified. However, such high levels of light output in a short time window ($\sim 1~\mu s$) can still cause the PMTs to saturate. This issue may be further mitigated by reading out the PMT signals from one of the last dynodes instead from the anode, or by optimizing the tapering ratio of the PMT voltage divider. It has been demonstrated that such saturation effects can also be partially corrected in analysis by focusing on the rising edges of the pulses [29]. Saturation leads to an additional uncertainty in the number of photons detected in the affected channel, which could degrade the performance of the SS/MS reconstruction. The following study ignores such photosensor saturation effects, assuming that either software or hardware solutions can be found to mitigate their impact in the experiment.
Sec. 2.3 estimates a theoretical limit of 7.3 mm for the benchmark MS2 rejection resolution with $1.2\times 10^{6}$ detected photons, assuming the use of 2" PMTs and a 1:9 photon splitting between two vertices. This benchmark resolution corresponds to a rejection of 84% MS2 backgrounds separated by this distance while accepting 50% SS signals. The background rejection power deteriorates for events with smaller separations between vertices or with more imbalanced energy distribution. Full calculations of the $0\nu\beta\beta$ signal acceptance and 214Bi background rejection power in our theoretical framework are not straightforward, so we use the MLE and CNN classifier to approximate the TPC’s capability in this application. As demonstrated in Sec. 3, these methods can approach the theoretical CRLB limits in idealized scenarios. Similar to earlier studies, we use an analytical light response function for both sensor hit pattern simulation and reconstruction, and thus ignore uncertainties beyond statistical fluctuations.
@@ 4.2 Experiment optimization
The CRLB derived from the FI matrix is an estimate of the best reconstruction accuracy for parameters of interest with an unbiased estimator. On the one hand, this limit may not be achieved in experiments, especially when simplifications are made in modeling the system behavior (forward model) during sensitivity predictions. Under favorable conditions, however, the limits could be reasonably approached; for example, the SS position resolution achieved by the LUX experiment is within 20% from the predictions made in Sec. 3.1.1. Further, the comparison of performance for different design options will remain valid if they share similar uncertainty sources. As a result, even when the predicted performance limit could not be reached, FI remains a useful optimization tool.
@@
Figure 17: Predicted SS (left) and MS2 (right) reconstruction resolutions for 2" photosensors (red lines) and 6 mm photosensors (blue lines) with different distances between the electroluminescence region and the photosensor array (denoted as $h$): 6 cm (solid lines), 1 cm (dashed lines), 0.5 cm (dotted lines). For MS2 events a 1:1 energy splitting between the two sites is assumed.
In our TPC model, the separation between MS event vertices is governed by physical laws and cannot be increased, while the observed spread of the light response function in a TPC can be strongly shaped by choices of detector configuration. Smaller sensor pixels can lead to sharper light response functions and better position resolution, as shown in Figs. 2 and 6. For sensor sizes comparable to, or smaller than the distance between the photon source and the sensor plane, the gain from reducing sensor size diminishes. Therefore, the benefits of small sensors can only be realized by choosing a small physical separation between the light source and the sensor plane. This is demonstrated in Fig. 17, where we compare the predicted SS (left) and MS2 (right, assuming 1:1 photon splitting between two vertices) resolutions for varying distances between the electroluminescence region and the photosensor plane, assuming the use of 2" sensors (red lines) and 6 mm sensors (blue lines). Compared to the baseline model of 1 cm sensor separation (dashed lines), a 0.5 cm separation  [^4] For the 0.5 cm sensor separation study, the electroluminescence region thickness is also reduced to 0.5 cm. (dotted lines) improves the resolution for 6 mm sensors by approximately a factor of 2; the results for 2" sensors remain largely unchanged because the photon distribution is dominated by the large size of the sensor in these two cases. When the sensor separation from the electroluminescence region is increased from 1 cm to 6 cm (solid lines), which was the configuration of the LUX detector, both SS and MS resolutions substantially deteriorate. In this configuration, the 2" sensor results and the 6 mm sensor results are nearly identical, because the photon distributions in both scenarios are dominated by the large light source-sensor separation.